An Overview of Radial Basis Functions

by

Greg Fasshauer

\begin{matrix}c \\ ◯ \end{matrix} August 8, 1996

REFERENCES

There are more references in [Hardy ’90] and in [Golberg & Chen ’96]. Those in [Hardy ’90] not mentioned below are mainly concerned with special applications of multiquadrics, those in [Golberg & Chen ’96] not covered below are from the application of radial basis functions in the dual reciprocity method for boundary element methods, and from the area of kriging in geostatistics.

Research Articles

1. Acosta, F. M., Radial basis functions and related models, Signal Proc. 45 (1995), 37–58.

2. Arad, N., N. Dyn, and D. Reisfeld, Image warping by radial basis functions: applications to facial expressions, CVGIP: Graphical models and image processing 56 (1994), 161–172.

3. Askey, R, Radial characteristic functions, TSR #1262, University of Wisconsin-Madison, 1973.

4. Ball, K., Eigenvalues of Euclidean distance matrices, J. Approx. Theory 68 (1992), 74–82.

5. Ball, K., N. Sivakumar, and J. D. Ward, On the sensitivity of radial basis interpolation to minimal data separation distance, Constr. Approx. 8 (1992), 401–426.

6. Barnes, R. J., A modified conjugate gradient algorithm for scattered data interpolation using radial basis functions, J. Appl. Sci. Comput. 1 (1994), 227–238.

7. Bates, D., M. Lindstrom, G. Wahba, and B. Yandell, GCVPACK - routines for generalized cross validation, Commun. Statist. B - Simulation and Computation 16 (1987), 263–297.

8. Bates, D. and G. Wahba, Computational methods for generalized cross validation with large data sets, in Treatment of Integral Equations by Numerical Methods, C. T. H. Baker, and G. F. Miller (eds), Academic Press, New York, 1982, 283–296.

9. Baxter, B. J. C., Conditionally positive functions and p-norm distance matrices, Constr. Approx. 7 (1991), 427–440.

10. Baxter, B. J. C., Norm estimates for inverses of distance matrices, in Mathematical Methods in Computer Aided Geometric Design II, T. Lyche and L. Schumaker (eds), Academic Press, New York, 1992, 9–18.

11. Baxter, B. J. C., The asymptotic cardinal function of the multiquadric ϕ(r) = (r^{2} + c^{2})^{1/2} as c → ∞, Comput. Math. Appl. 24 (1992), 1–6.

12. Baxter, B. J. C., The interpolation theory of radial basis functions, dissertation, University of Cambridge, 1992.

13. Baxter, B. J. C., N. Sivakumar, and J. D. Ward, Regarding the p-norms of radial basis interpolation matrices, Constr. Approx. 10 (1994), 451–468.

14. Beatson, R. K. and W. A. Light, Quasi-interpolation in the absence of polynomial reproduction, in Numerical Methods in Approximation Theory, ISNM 105, D. Braess, L.L. Schumaker (ed), Birkhäuser, Basel, 1992, 21–39.

15. Beatson, R. K. and W. A. Light, Quasi-interpolation with thin plate splines on a square, Constr. Approx. 9 (1993), 407–433.

16. Beatson, R. K. and G. N. Newsam, Fast evaluation of radial basis functions: I, Comput. Math. Appl. 24 (1992), 7–19.

17. Beatson, R. K. and M. J. D. Powell, Univariate multiquadric approximation: quasi-interpolation to scattered data, Constr. Approx. 8 (1992), 275–288.

18. Beatson, R. K. and M. J. D. Powell, Univariate interpolation on a regular finite grid by a multiquadric plus a linear polynomial, J. Inst. Math. Applics. 12 (1992), 107–133.

19. Beatson, R. K. and M. J. D. Powell, An iterative method for thin plate spline interpolation that employs approximations to Lagrange functions, in Numerical Analysis 1993 (Dundee, 1993), xx (ed), Pitman Res. Notes Math. Ser., 303, Longman Sci. Tech., Harlow, 1994, 17–39.

20. Binev, P. and K. Jetter, Estimating the condition number for multivariate interpolation problems, in Numerical Methods in Approximation Theory, ISNM 105, D. Braess, L.L. Schumaker (ed), Birkhäuser, Basel, 1992, 41–52.

21. Bingham, N. H., Positive definite functions on spheres, Proc. Camb. Phil. Soc. 73 (1973), 145–156.

22. Bishop, C., Improving the generalization properties of radial basis function neural networks, Neural Computation 3 (1991), 579–588.

23. Bochner, S., Vorlesungen über Fouriersche Integrale, Akademische Verlagsgesellschaft, Leipzig, 1932.

24. Bochner, S., Monotone Funktionen, Stieltjes Integrale und harmonische Analyse, Math. Ann. 108 (1933), 378–410.

25. Bochner, S., Hilbert distances and positive definite functions, Ann. of Math. 42 (1941), 647–656.

26. de Boor, C., R. A. DeVore, and A. Ron, Approximation from shift-invariant subspaces of L_{2}(IR^{d}), Trans. Amer. Math. Soc. 341 (1994), 787–806.

27. de Boor, C., R. A. DeVore, and A. Ron, The structure of finitely generated shift-invariant spaces in L_{2}(IR^{d}), J. Funct. Anal. 119 (1994), 37–78.

28. de Boor, C. and A. Ron, Fourier analysis of the approximation power of principal shift-invariant spaces, Constr. Approx. 8 (1992), 427–462.

29. Bos, L. P. and K. Šalkauskas, On the matrix [|x_{i} − x_{j}|^{3}] and the cubic spline continuity equations, J. Approx. Theory 51 (1987), 81–88.

30. Broomhead, D. S. and D. Lowe, Multivariate functional interpolation and adaptive networks, Complex Systems 2 (1988), 321–355.

31. Brown, A. L., Uniform approximation by radial basis functions, Appendix B to Radial basis functions in 1990, in Advances in Numerical Analysis II: Wavelets, Subdivision, and Radial Basis Functions, W. Light (ed), Oxford University Press, Oxford, 1992, 203–206.

32. Buckley, M. J., Fast computation of a discretized thin-plate smoothing spline for image data, Biometrika 81 (1994), 247–258.

33. Buhmann, M. D., Convergence of univariate quasi-interpolation using multiquadrics, IMA J. Numer. Anal. 8 (1988), 365–383.

34. Buhmann, M. D., Multivariate interpolation using radial basis functions, dissertation, University of Cambridge, 1989.

35. Buhmann, M. D., Cardinal interpolation with radial basis functions: an integral approach, in Multivariate Approximation Theory IV, ISNM 90, C. Chui, W. Schempp, and K. Zeller (eds), Birkhäuser Verlag, Basel, 1989, 41–64.

36. Buhmann, M. D., Multivariate interpolation in odd-dimensional Euclidean spaces using multiquadrics, Constr. Approx. 6 (1990), 21–34.

37. Buhmann, M. D., Multivariate cardinal interpolation with radial basis functions, Constr. Approx. 6 (1990), 225–255.

38. Buhmann, M. D., New developments in the theory of radial basis function interpolation, in Multivariate Approximation: From CAGD to Wavelets, Kurt Jetter and Florencio Utreras (eds), World Scientific Publishing, Singapore, 1993, 35–75.

39. Buhmann, M. D., On quasi-interpolation with radial basis functions, J. Approx. Theory 72 (1993), 103–130.

40. Buhmann, M. D., Discrete least squares approximation and prewavelets from radial function spaces, Proc. Camb. Phil. Soc. 114 (1993), 533–558.

41. Buhmann, M. D., Multiquadric prewavelets on nonequally spaced knots in one dimension, Math. Comp. 64 (1995), 1611–1625.

42. Buhmann, M. D., Pre-wavelets on scattered knots and from radial function spaces: a review, in The Mathematics of Surfaces VI, R R. Martin (ed), Clarendon Press, Oxford, 1995, xxx–xxx.

43. Buhmann, M. D., Radial functions on compact support, Research Rep. 95-13, ETH Zürich, 1995.

44. Buhmann, M. D. and C. K. Chui, A note on the local stability of translates of radial basis functions, J. Approx. Theory 74 (1993), 36–40.

45. Buhmann, M. D., F. Derrien, and A. Le Méhauté, Spectral properties and knot removal for interpolation by pure radial sums, in Mathematical Methods for Curves and Surfaces, M. Dæhlen, T. Lyche, and L. Schumaker (eds), Vanderbilt University Press, Nashville, 1995, 55–62.

46. Buhmann, M. D. and N. Dyn, Error estimates for multiquadric interpolation, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 51–58.

47. Buhmann, M. D. and N. Dyn, Spectral convergence of multiquadric interpolation, Proc. Edinburgh Math. Soc. 36 (1993), 319–333.

48. Buhmann, M. D., N. Dyn, and D. Levin, On quasi-interpolation with radial basis functions on non-regular grids, Constr. Approx. 11 (1995), 239–254.

49. Buhmann, M. D. and A. Le Méhauté, Knot removal with radial basis function interpolantion, C. R. Acad. Sci. Paris Sér. I Math. 320 (1995), 501–506.

50. Buhmann, M. D. and C. A. Micchelli, Completely monotonic functions for cardinal interpolation, in Approximation Theory VI, C. Chui, L. Schumaker, and J. Ward (eds), Academic Press, New York, 1989, 1–4.

51. Buhmann, M. D. and C. A. Micchelli, Multiply monotone functions for cardinal interpolation, Adv. in Appl. Math. 12 (1991), 358–386.

52. Buhmann, M. D. and C. A. Micchelli, Multiquadric interpolation improved, Comput. Math. Appl. 24 (1992), 21–25.

53. Buhmann, M. D. and C. A. Micchelli, On radial basis approximation on periodic grids, Proc. Camb. Phil. Soc. 112 (1992), 317–334.

54. Buhmann, M. D. and P. Pinkus, On a recovery problem, Research Rep. 95-16, ETH Zürich, 1995.

55. Buhmann, M. D. and M. J. D. Powell, Radial basis function interpolation on an infinite regular grid, in Algorithms for Approximation II, M. G. Cox and J. C. Mason (eds), Chapman & Hall, London, 1990, 146–169.

56. Buhmann, M. D. and A. Ron, Radial basis functions: L^{p}-approximation orders with scattered centres, in Wavelets, Images, and Surface Fitting, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), A. K. Peters, Wellesley MA, 1994, 93–112.

57. Caiti, A., G. Magenes, T. Panisini, and R. Simpson, Smooth approximation by radial basis functions: three case studies, J. Appl. Sc. Comp. 1 (1994), 88–113.

58. Carlson, R. E. and T. A. Foley, The parameter R^{2} in multiquadric interpolation, Comput. Math. Appl. 21 (1991), 29–42.

59. Carlson, R. E. and T. A. Foley, Interpolation of track data with radial basis methods, Comput. Math. Appl. 24 (1992), 27–34.

60. Carlson, R. E. and B. K. Natarajan, Sparse approximate multiquadric interpolation, Comput. Math. Appl. 27 (1994), 99–108.

61. Chan, A. K., C. K. Chui, and L. T. Guan, Radial basis function approach to interpolation of large reflecting surfaces, in Curves and Surfaces in Computer Vision and Graphics, L. A. Ferrari, and R. J. P. de Figueiredo (eds), SPIE, Vol. 1251, 1990, xxx–xxx.

62. Chang, K.-F., Strictly positive definite functions, Feng Chia University, 1994.

63. Chang, K.-F., Strictly positive definite functions II, Feng Chia University, 1996.

64. Chen, S., Radial basis functions for signal prediction and system modelling, J. Appl. Sci. Comput. 1 (1994), 172–207.

65. Chen, S., S. A. Billings, and P. M. Grant, Recursive hybrid algorithm for nonlinear system identification using radial basis function networks, Int. J. Control 55 (1992), 1051–1070.

66. Chen, T. and H. Chen, Approximation capability to functions of several variables, nonlinear functionals and operators by radial basis function neural networks, IEEE Trans. Neural Networks 6 (1995), 904–910.

67. Chen, T. and H. Chen, Denseness of radial-basis functions in L^{2}(IR^{n}) and its applications in neural networks, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 137–144.

68. Chen, S., G. J. Gibson, C. F. N. Cowan, and P. M. Grant, Reconstruction of binary signals using an adaptive radial-basis-function equalizer, Signal Processing 22 (1991), 77–93.

69. Cheney, E. W., Approximation by functions of nonclassical form, in Approximation Theory, Spline Functions and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1992, 1–18.

70. Cheney, E. W., Approximation and interpolation on spheres, in Approximation Theory, Wavelets and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1995, 47–53.

71. Cheney, E. W., Approximation using positive definite functions, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 145–168.

72. Cheney, E. W., W. A. Light, and Y. Xu, On kernels and approximation orders, in Proc. Sixth Southeastern Approximation Theory Conference, George Anastassiou (ed), Lecture Notes in Pure and Applied Mathematics, Vol. 138, 1992, 227–242.

73. Cheney, E. W. and Y. Xu, A set of research problems in approximation theory, in Topics in Polynomials of One and Several Variables and Their Applications, T. M. Rassias, H. M. Srivastava, and A. Yanushauskas (eds), World Scientific Publishers, London, 1993, 109–123.

74. Cheng, C. C. and Y. F. Zheng, Thin plate spline surface approximation using Coons’ patches, Comput. Aided Geom. Design 11 (1994), 269–287.

75. Chui, C. K., J. Stöckler, and J. D. Ward, Analytic wavelets generated by radial functions, Adv. Comput. Math. 5 (1996), 95–123.

76. Chui, C. K., J. D. Ward, and K. Jetter, Cardinal interpolation with differences of tempered functions, Comput. Math. Appl. 24 (1992), 35–48.

77. Delvos, F. J., Convergence of interpolation by translation, Colloq. Math. Soc. J. Bolya 49 (1985), 273–287.

78. Delvos, F. J., Periodic interpolation on uniform meshes, J. Approx. Theory 51 (1987), 71–80.

79. Dix, J. G. and R. D. Ogden, An interpolation scheme with radial basis in Sobolev spaces H^{s}(IR^{n}), Rocky Mountain J. Math. 24 (1994), 1319–1337.

80. Dubal, M. R., Construction of three-dimensional black-hole initial data via multiquadrics, Phys. Rev. D 45 (1992), 1178–1187.

81. Dubal, M. R., Domain decomposition and local refinement for multiquadric approximations. I: Second-order equations in one-dimension, J. Appl. Sc. Comp. 1 (1994), 146–171.

82. Dubal, M. R., S. R. Oliveira, and R. A. Matzner, Solution of elliptic equations in numerical relativity using multiquadrics, in Approaches to Numerical Relativity, R. d’Inverno (ed), Cambridge University Press, Cambridge, 1992, 265–280.

83. Duchon, J., Interpolation des fonctions de deux variables suivant le principe de la flexion des plaques minces, Rev. Française Automat. Informat. Rech. Opér., Anal. Numer. 10 (1976), 5–12.

84. Duchon, J., Splines minimizing rotation-invariant semi-norms in Sobolev spaces, in Constructive Theory of Functions of Several Variables, Oberwolfach 1976, W. Schempp and K. Zeller (eds), Springer Lecture Notes in Math. 571, Springer-Verlag, Berlin, 1977, 85–100.

85. Duchon, J., Sur l’erreur d’interpolation des fonctions de plusieurs variables par les D^{m}-splines, Rev. Française Automat. Informat. Rech. Opér., Anal. Numer. 12 (1978), 325–334.

86. Duchon, J., Fontions splines homogènes à plusiers variables, dissertation, Université de Grenoble, 1980.

87. Dyn, N., Interpolation of scattered data by radial functions, in Topics in Multivariate Approximation, C. K. Chui, L. L. Schumaker, and F. Utreras (eds), Academic Press, New York, 1987, 47–61.

88. Dyn, N., Interpolation and approximation by radial and related functions, in Approximation Theory VI, C. Chui, L. Schumaker, and J. Ward (eds), Academic Press, New York, 1989, 211–234.

89. Dyn, N., T. Goodman, and C. A. Micchelli, Positive powers of certain conditionally negative definite matrices, Indaga. Math. Proc. Ser.A. 89 (1986), 163–178.

90. Dyn, N., I. R. H. Jackson, D. Levin, and A. Ron, On multivariate approximation by the integer translates of a basis function, Israel J. Math. 78 (1992), 95–130.

91. Dyn, N. and D. Levin, Bell shaped basis functions for surface fitting, in Approximation Theory and Applications, Z. Ziegler (ed), Academic Press, New York, 1981, 113–129.

92. Dyn, N. and D. Levin, Construction of surface spline interpolants of scattered data over finite domains, Rev. Française Automat. Informat. Rech. Opér., Anal. Numer. 16 (1982), 201–209.

93. Dyn, N. and D. Levin, Iterative solution of systems originating from integral equations and surface interpolation, SIAM J. Numer. Anal. 20 (1983), 377–390.

94. Dyn, N., D. Levin, and S. Rippa, Surface interpolation and smoothing by “thin plate” splines, in Approximation Theory IV, C. Chui, L. Schumaker, and J. Ward (eds), Academic Press, New York, 1983, 445–449.

95. Dyn, N., D. Levin, and S. Rippa, Numerical procedures for surface fitting of scattered data by radial functions, SIAM J. Sci. Statist. Comput. 7 (1986), 639–659.

96. Dyn, N., W. A. Light, and E. W. Cheney, Interpolation by piecewise-linear radial basis functions, I, J. Approx. Theory 59 (1989), 202–223.

97. Dyn, N. and C. A. Micchelli, Interpolation by sums of radial functions, Numer. Math. 58 (1990), 1–9.

98. Dyn, N. and A. Ron, Radial basis function approximation: from gridded centers to scattered centers, Proc. London Math. Soc. 71 (1995), 76–108.

99. Dyn, N. and A. Ron, Multiresolution analysis by infinitely differentiable compactly supported functions, Appl. Comput. Harmon. Anal. 2 (1995), 15–20.

100. Fasshauer, G. E., Scattered data interpolation with radial basis functions on the sphere, Vanderbilt University, 1994.

101. Fasshauer, G. E., On the density of certain classes of radial basis functions on spheres, Vanderbilt University, 1994.

102. Fasshauer, G. E., Adaptive least squares fitting with radial basis functions on the sphere, in Mathematical Methods for Curves and Surfaces, M. Dæhlen, T.

Lyche, and L. Schumaker (eds), Vanderbilt University Press, Nashville, 1995, 141–150.

103. Fasshauer, G. E., Hermite interpolation with radial basis functions on spheres, submitted for publication, 1995.

104. Fasshauer, G. E., Radial Basis Functions on Spheres, dissertation, Vanderbilt University, 1995.

105. Fasshauer, G. E., Solving partial differential equations by collocation with radial basis functions, Northwestern University, 1996.

106. Floater, M. S. and A. Iske, Thinning and approximation of large sets of scattered data, in Advanced Topics in Multivariate Approximation, F. Fontanella, K. Jetter, and P.-J. Laurent (eds), World Scientific Publishing, Singapore, 1996, xxx–xxx.

107. Floater, M. S. and A. Iske, Multistep scattered data interpolation using compactly supported radial basis functions, J. Comput. Applied math., to appear.

108. Foley, T. A., Interpolation and approximation of 3-D and 4-D scattered data, Comput. Math. Appl. 13 (1987), 711–740.

109. Foley, T. A., Interpolation of scattered data on a spherical domain, in Algorithms for Approximation II, M. G. Cox and J. C. Mason (eds), Chapman & Hall, London, 1990, 303–310;.

110. Foley, T. A., The map and blend scattered data interpolant on the sphere, Comput. Math. Appl. 24 (1992), 49–60.

111. Foley, T. A., Near optimal parameter selection for multiquadric interpolation, J. Appl. Sc. Comp. 1 (1994), 54–69.

112. Foley, T. A., S. Dayanand, and D. Zeckzer, Localized radial basis methods using rational triangle patches, in Geometric Modelling – Dagstuhl 1993, H. Hagen, G. Farin and H. Noltemeier (eds), Springer Verlag, Berlin, 1995, 163– 176.

113. Franke, R., Scattered data interpolation: tests of some methods, Math. Comp. 48 (1982), 181–200.

114. Franke, R., Smooth interpolation of scattered data by local thin plate splines, Comput. Math. Appl. 8 (1982), 273–281.

115. Franke, R., Thin plate splines with tension, Comput. Aided Geom. Design 2 (1985), 87–95.

116. Franke, R., Recent advances in the approximation of surfaces from scattered data, in Topics in Multivariate Approximation, C. K. Chui, L. L. Schumaker, and F. Utreras (eds), Academic Press, New York, 1987, 79–98.

117. Franke, R., Approximation of scattered data for meteorological applications, in Multivariate Approximation and Interpolation, ISNM 94, W. Haussman and K. Jetter (eds), Birkhäuser, Basel, 1990, 107–120.

118. Franke, R., Paper presented at SIAM Geometric Modeling Conference in Nashville Nov. 1995, SIAM, 1995.

119. Franke, R., H. Hagen, and G. M. Nielson, Least squares surface approximation to scattered data using multiquadric functions, Adv. in Comput. Math. 2 (1994), 81-99.

120. Franke, R., H. Hagen, and G. M. Nielson, Repeated knots in least squares multiquadric functions, in Geometric Modelling – Dagstuhl 1993, H. Hagen, G. Farin and H. Noltemeier (eds), Springer Verlag, Berlin, 1995, 177-187.

121. Franke, R. and G. M. Nielson, Scattered data interpolation and applications: A tutorial and survey, in Geometric Modeling, Methods and Applications, H. Hagen and D. Roller (eds), Springer Verlag, Berlin, 1991, 131–160.

122. Franke, R. and K. Šalkauskas, Localization of multivariate interpolation and smoothing methods, Paper presented at SIAM Geometric Modeling Conference, Nashville, 1995.

123. Freeden, W., Spherical spline interpoaltion – basic theory and computational aspects, J. Comput. Appl. Math. 11 (1984), 367–375.

124. Freeman, J. A. and D. Saad, Learning and generalization in radial basis function networks, Neural Computation 7 (1995), 1000-1020.

125. Galperin, E. A., E. J. Kansa, A. Makroglou, and S. A. Nelson, Numerical solutions of weakly singular Volterra equations using global optimization techniques and radial basis functions, University of Quebec, 1996.

126. Galperin, E. A. and Q. Zheng, Solution and control of PDE via global optimization methods, Comput. Math. Appl. 25 (1993), 103–118.

127. Gibson, R. and G. Stone, Computation of thin plate splines, SIAM J. Sci. Statist. Comput. 12 (1991), 1304–1313.

128. Girosi, F., On some extensions of radial basis functions and their applications in artificial intelligence, Comput. Math. Appl. 24 (1992), 61–80.

129. Girosi, F. and G. Anzellotti, Rates of convergence for radial bais functions and neural networks, in Artificial Networks for Speech and Vision, R. J. Mammone (ed), Chapman & Hall, London, 1993, xxx–xxx.

130. Girosi, F., M. Jones, and T. Poggio, Regularization theory and neural network architectures, MIT AI memo 1430, 1993.

131. Golberg, M. A., A note on the sparse representation of discrete integral operators, Appl. Math. Comp. 70 (1995), 97–118.

132. Golberg, M. A., Recent developments in the numerical evaluation of particular solutions in the boudary element method, Appl. Math. Comp. 75 (1996), 91– 101.

133. Golberg, M. A. and C. S. Chen, The theory of radial basis functions applied to the BEM for inhomogeneous partial differential equations, Boundary Elements Comm. 5 (1994), 57–61.

134. Golberg, M. A. and C. S. Chen, A bibliography on radial basis function approximation, UNLV, 1996.

135. Golberg, M. A., C. S. Chen, and S. R. Karur, Improved multiquadric approximation for partial differential equations, University of Nevada, 1995.

136. Gu, C. and G. Wahba, Semiparametric analysis of variance with tensor product thin plate splines, J. Roy. Statist. Soc. Ser. B 55 (1993), 353–368.

137. Guo, K., S. Hu, and X. Sun, Conditionally positive definite functions and Laplace-Stieltjes integrals, J. Approx. Theory 74 (1993), 249–265.

138. Guo, K., S. Hu, and X. Sun, Cardinal interpolation using linear combinations of translates of conditionally positive definite functions, Numer. Func. Anal. Optim. 14 (1993), 371–381.

139. Guo, K. and X. Sun, Scattered data interpolation by linear combinations of translates of conditionally positive definite functions, Numer. Func. Anal. Optim. 12 (1991), 137–152.

140. Gutzmer, T., Interpolation by positive definite functions on locally compact groups with application to SO(3), Results Math. 29 (1996), 69–77.

141. Hagan, R. E., Using multiquadric and general finite differences in ground water modeling, dissertation, Utah State University, 19xx.

142. Hagan, R. E., Multiquadrics in engineering modeling, in Proceedings ASCE National Conference on Irrigation and Drainage Engineering, Park City, Utah, 1993.

143. Hagan, R. E. and E. J. Kansa, Studies of the R parameter in the multiquadric function applied to ground water pumping, J. Appl. Sci. and 1 (1994), 266–281.

144. Halgamuge, S. K., W. Poechmueller, and M. Glesner, An alternative approach for generation of membership functions and fuzzy rules based on radial and cubic basis function networks, Internat. J. Approx. Reason. 12 (1995), 279– 298.

145. Halton, E. J. and W. A. Light, On local and controlled approximation order, J. Approx. Theory 72 (1993), 268–277.

146. Handscomb, D., Local recovery of a solenoidal vector field by an extension of the thin-plate spline technique, Algorithms for approximation, III (Oxford, 1992), Numer. Algorithms 5, 1993, no. 1-4, 121–129.

147. Harder, R. L. and R. N. Desmarais, Interpolation using surface splines, J. Aircraft 9 (1972), 189–191.

148. Hardy, R. L., Multiquadric equations of topography and other irregular surfaces, J. Geophys. Res. 76 (1971), 1905–1915.

149. Hardy, R. L., Geodetic applications of multiquadric analysis, AVN Allg. Vermess. Nachr. 79 (1972), 389–406.

150. Hardy, R. L., Research results in the application of multiquadric equations to surveying and mapping problems, Survg. Mapp. 35 (1975), 321–332.

151. Hardy, R. L., Least squares prediction, Photogramm. Eng. and Remote Sensing 43 (1977), 475–492.

152. Hardy, R. L., Theory and applications of the multiquadric-biharmonic method, Comput. Math. Appl. 19 (1990), 163–208.

153. Hardy, R. L., Comments on the discovery and evolution of the multiquadric method, Comput. Math. Appl. 24 (1992), xi–xii.

154. Hardy, R. L., A contribution of the multiquadric method: interpolation of potential inside the earth, Comput. Math. Appl. 24 (1992), 81–97.

155. Hardy, R. L. and W. M. Göpfert, Least squares prediction of gravity anomalies, geodial undulations, and deflections of the vertical with multiquadric harmonic functions, Geophys. Res. Letters 2 (1975), 423–426.

156. Hardy, R. L. and S. A. Nelson, A multiquadric-biharmonic representation and approximation of disturbing potential, Geophys. Res. Letters 13 (1986), 18–21.

157. Hardy, R. L. and S. Siranyone, The multiquadric-biharmonic method of a three dimensional mapping inside ore deposits, in Proceedings of Multinational Conference on Mine Planning and Design, University of Kentucky, 1989.

158. Hutchinson, M. F., The application of thin plate smoothing splines to continentwide data assimilation, in BMRC Research Report No.27, Data Assimilation Systems, J. D. Jasper (ed), Bureau of Meteorology, Melbourne, 1991, 104–113.

159. Hutchinson, M. F., On thin plate splines and kriging, in Computing and Science in Statistics 25, M. E. Tarter, and M. D. Lock (eds), Interface Foundation of North America, University of California, Berkeley, 1993, 55–62.

160. Hutchinson, M. F., Interpolating mean rainfall using thin plate smoothing splines, Int. J. GIS 9 (1995), 305–403.

161. Iske, A., Charakterisierung bedingt positiv definiter Funktionen für multivariate Interpolationsmethoden mit radial Basisfunktionen, dissertation, Universität Göttingen, 1994.

162. Iske, A., Reconstruction of functions from generalized Hermite-Birkhoff data, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 257–264.

163. Iske, A. and T. Sonar, On the structure of function spaces in optimal recovery of point functionals for ENO-schemes by radial basis functions, Universität Göttingen, 1995.

164. Jackson, I. R. H., Approximation to continuous functions using radial basis functions, in The Mathematics of Surfaces II, R R. Martin (ed), Clarendon Press, Oxford, 1987, 115–135.

165. Jackson, I. R. H., Radial basis function methods from multivariate approximation, dissertation, University of Cambridge, 1988.

166. Jackson, I. R. H., Convergence properties of radial basis functions, Constr. Approx. 4 (1988), 243–264.

167. Jackson, I. R. H., An order of convergence for some radial basis functions, IMA J. Numer. Anal. 9 (1989), 567–587.

168. Jackson, I. R. H., Radial basis functions – a survey and new results, in Mathematics of Surfaces III, D. C. Handscomb (ed), Clarendon Press, Oxford, 1989, 115–133.

169. Jekeli, C., Hardy’s multiquadric-biharmonic method for gravity field predictions, Comput. Math. Appl. 28 (1994), 43–46.

170. Jetter, K., Multivariate approximation from the cardinal interpolation point of view, in Approximation Theory VII, E. W. Cheney, C. Chui, and L. Schumaker (eds), Academic Press, New York, 1992, 131–161.

171. Jetter, K., Riesz bounds in scattered data interpolation and L_{2}-approximation, in Multivariate Approximation: From CAGD to Wavelets, Kurt Jetter and Florencio Utreras (eds), World Scientific Publishing, Singapore, 1993, 167– 177.

172. Jetter, K., Conditionally lower Riesz bounds in scattered data interpolation, in Wavelets, Images, and Surface Fitting, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), A. K. Peters, Wellesley MA, 1994, 295–302.

173. Jetter, K. and J. Stöckler, Algorithms for cardinal interpolation using box splines and radial basis functions, Numer. Math. 60 (1991), 97–114.

174. Jia, R. Q. and J. J. Lei, Approximation by multiinteger translates of functions having global support, J. Approx. Theory 72 (1991), 2–23.

175. Jia, R. Q. and J. J. Lei, A new version of the Strang-Fix conditions, J. Approx. Theory 74 (1993), 221-225.

176. Kansa, E. J., Application of Hardy’s multiquadric interpolation to hydrodynamics, Proc. 1986 Simul. Conf., Vol. 4, (1986), 111–117.

177. Kansa, E. J., Multiquadrics – A scattered data approximation scheme with applications to computational fluid-dynamics – I: Surface approximations and partial derivative estimates, Comput. Math. Appl. 19 (1990), 127–145.

178. Kansa, E. J., Multiquadrics – A scattered data approximation scheme with applications to computational fluid-dynamics – II: Solutions to parabolic, hyperbolic and elliptic partial differential equations, Comput. Math. Appl. 19 (1990), 147–161.

179. Kansa, E. J., A strictly conservative spatial approximation scheme for the governing engineering and physics equations over irregular regions and inhomogeneous scattered nodes, Comput. Math. Appl. 24 (1992), 169–190.

180. Kansa, E. J. and R. E. Carlson, Improved accuracy of multiquadric interpolation using variable shape parameters, Comput. Math. Appl. 24 (1992), 99–120.

181. Kansa, E. J. and R. E. Carlson, Radial basis functions: a class of grid free scattered data approximations, Computational Fluid Dynamics J. 3 (1995), 479–496.

182. Karur, S. R. and P. A. Ramachandran, Radial basis function approximation in the dual reciprocity method, Math. Comput. Modell. 20 (1994), 59–70.

183. Karur, S. R. and P. A. Ramachandran, Augmented thin plate spline approximation in DRM, Boundary Elements Comm. 6 (1995), 55–58.

184. Kent, J. T. and K. V. Mardia, The link between kriging and thin-plate splines, Probability, statistics and optimisation, Wiley Ser. Probab. Math. Statist., Wiley, Chichester, 1994, 325–339.

185. Kurdila, A. J., F. J. Narcowich, and J. D. Ward, Persistency of excitation in identification using radial basis function approximants, SIAM J. Control Optim. 33 (1995), 625–642.

186. Lee, D. and J. H. Shiau, Thin plate splines with discontinuities and fast algorithms for their computation, SIAM J. Scient. Computing 15 (1994), 1311-1330.

187. Le Méhauté, A., Knot removal for scattered data, in Approximation Theory, Wavelets and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1995, 197–213.

188. Leonard, J. A., M. A. Kramer, and L. W. Ungar, Using radial basis functions to approximate a function and its error bounds, IEEE Trans. on Neural Networks 3 (1992), 624.

189. Levesley, J., Convolution kernels for approximation by radial basis functions, in Wavelets, Images, and Surface Fitting, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), A. K. Peters, Wellesley MA, 1994, 343–350.

190. Levesley, J., Pointwise estimates for multivariate interpolation using conditionally positive definite functions, in Approximation Theory, Wavelets and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1995, 381– 401.

191. Levesley, J., Convolution kernels based on thin plate splines, Numer. Algorithms 10 (1995), 401–419.

192. Levesley, J. and A. K. Kushpel, Interpolation on compact Abelian groups using generalised sk-splies, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 317–324.

193. Levesley, J. and A. K. Kushpel, Generalised sk-spline interpolation on compact Abelian groups, University of Leicester, 1995.

194. Levesley, J. and M. Roach, Quasi-interpolation on compact domains, in Approximation Theory, Wavelets and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1995, 557–566.

195. Levesley, J. and X. Sun, Scattered Hermite interpolation by ridge functions, Numer. Func. Anal. Optim. 16 (1995), 989–1001.

196. Levesley, J., Y. Xu, W. Light, and W. Cheney, Convolution operators for radial basis approximation, SIAM J. Math. Anal. 27 (1996), 286–304.

197. Light, W. A., Recent developments in the Strang-Fix theory for approximation orders, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 285–292.

198. Light, W. A., Some aspects of radial basis function approximation, in Approximation Theory, Spline Functions and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1992, 163–190.

199. Light, W. A., Using radial functions on compact domains, in Wavelets, Images, and Surface Fitting, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), A. K. Peters, Wellesley MA, 1994, 351–370.

200. Light, W. A. and E. W. Cheney, Interpolation by piecewise-linear radial basis functions, II, J. Approx. Theory 64 (1991), 38–54.

201. Light, W. A. and E. W. Cheney, Interpolation by periodic radial basis functions, J. Math. Anal. Appl. 168 (1992), 111–130.

202. Light, W. A. and E. W. Cheney, Quasi-interpolation with translates of a function having noncompact support, Constr. Approx. 8 (1992), 35–48.

203. Light, W. A. and H. Wayne, Error estimates for approximation by radial basis functions, in Approximation Theory, Wavelets and Applications, S. P. Singh (ed), Kluwer, Dordrecht, Netherlands, 1995, 215–246.

204. Locher, F., Interpolation on uniform meshes by the translates of one function and related attenuation factors, Math. Comp. 37 (1981), 403–416.

205. Louse, D. and R. Matthews, Shakespeare vs. Fletcher: a stylistic analysis by radial basis functions, Computers and Humanities 29 (1995), 449–xx.

206. Lowe, D. and A. McLachlan, Modelling of nonstationary processes using radial basis function networks, Neural Computing Research Group, Aston University, UK, 199x.

207. Luo, Z., The invertibility of interpolation functions and the fundamentality of their translates on spheres, preprint, 1994.

208. Luo, Z., Characterization and eigenvalue estimation of conditionally positive definite functions on spheres, preprint, 1994.

209. Madych, W. R., Cardinal interpolation with polyharmonic splines, in Multivariate Approximation Theory IV, ISNM 90, C. Chui, W. Schempp, and K. Zeller (eds), Birkhäuser Verlag, Basel, 1989, 241–248.

210. Madych, W. R., Polyharmonic splines, multiscale analysis, and entire functions, in Multivariate Approximation and Interpolation, ISNM 94, W. Haussman and K. Jetter (eds), Birkhäuser, Basel, 1990, 205–216.

211. Madych, W. R., Error estimates for interpolation by generalized splines, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 297–306.

212. Madych, W. R., Miscellaneous error bounds for multiquadric and related interpolators, Comput. Math. Appl. 24 (1992), 121–138.

213. Madych, W. R. and S. A. Nelson, Multivariate interpolation: a variational theory, manuscript, 1983.

214. Madych, W. R. and S. A. Nelson, Multivariate interpolation and conditionally positive definite functions, Approx. Theory Appl. 4 (1989), 77–89.

215. Madych, W. R. and S. A. Nelson, Error bounds for multiquadric interpolation, in Approximation Theory VI, C. Chui, L. Schumaker, and J. Ward (eds), Academic Press, New York, 1989, 413–416.

216. Madych, W. R. and S. A. Nelson, Multivariate interpolation and conditionally positive definite functions, II, Math. Comp. 54 (1990), 211–230.

217. Madych, W. R. and S. A. Nelson, Polyharmonic cardinal splines, J. Approx. Theory 60 (1990), 141–156.

218. Madych, W. R. and S. A. Nelson, Polyharmonic cardinal splines: A minimization property, J. Approx. Theory 63 (1990), 303–320.

219. Madych, W. R. and S. A. Nelson, Bounds on multivariate polynomials and exponential error estimates for multiquadric interpolation, J. Approx. Theory 70 (1992), 94–114.

220. Makroglu, A., Radial basis functions in the numerical solution of Fredholm and integro-differential equations, in Advances in Computer Methods for Partial Differential Equations – VII, R. Vichnevetsky, D. Knight, and G. Richter (eds), IMACS, New Brunswick, 1992, 478–484.

221. Makroglu, A., Radial basis functions in the numerical solution of nonlinear Volterra integral equations, J. Appl. Sci. Comput. 1 (1994), 33–53.

222. Makroglu, A., Multiquadric collocation methods in the numerical solution of Volterra integral and integro-differential equations, Mathematics of Computation 1943–1993, Proc. Sympos. Appl. Math., 48, Amer. Math. Soc., Providence, RI, 1994, 337–341.

223. Makroglu, A. and E. J. Kansa, Multiquadric collocation methods in the numerical solution of Volterra integral equations with weakly singular kernels, Lawrence Livermore National Laboratory, 1993.

224. McMahon, J., Knot selection for least squares approximation using thin plate splines, M.S. Thesis, Naval Postgraduate School, 1986.

225. McMahon, J. and R. Franke, Knot selection for least squares thin plate splines, SIAM J. Sci. Statist. Comput. 13 (1992), 484–498.

226. Meinguet, J., Multivariate interpolation at arbitrary points made simple, Z. Angew. Math. Phys. 30 (1979), 292–304.

227. Meinguet, J., An intrinsic approach to multivariate spline interpolation at arbitrary points, in Polynomial and Spline Approximations, N. B. Sahney (ed), Reidel, Dordrecht, 1979, 163–190.

228. Meinguet, J., Surface spline interpolation: basic theory and computational aspects, in Approximation Theory and Spline Functions, S. P. Singh, J. H. W. Burry, and B. Watson (eds), Reidel, Dordrecht, 1984, 127–142.

229. Menegatto, V. A., Interpolation on spherical domains, Analysis 14 (1994), 415–424.

230. Menegatto, V. A., Strictly positive definite kernels on the Hilbert sphere, Appl. Anal. 55 (1994), 91–101.

231. Menegatto, V. A., Interpolation on the complex Hilbert sphere, Approx. Theory Appl. 11 (1995), 1–9.

232. Menegatto, V. A., Strictly conditionally negative definite functions on the sphere, University of Texas Austin, 1992.

233. Menegatto, V. A., Interpolation on Spherical Spaces, dissertation, University of Texas Austin, 1992.

234. Menegatto, V. A., Strictly conditionally positive definite functions on the unit Hilbert sphere, University of Texas Austin, 1992.

235. Menegatto, V. A., Strictly positive definite kernels on the circle, Rocky Mountain J. Math. 25 (1995), 1149–1163;.

236. Menegatto, V. A., Interpolation on the complex Hilbert sphere using positive definite and conditionally negative definite functions, manuscript, 1995.

237. Mhaskar, H. N., Weighted polynomials, radial basis functions and potentials on locally compact spaces, Numer. Func. Anal. Optim. 11 (1990), 987–1017.

238. Mhaskar, H. N. and C. A. Micchelli, Approximation by superposition of sigmoidal and radial basis functions, Adv. in Appl. Math. 13 (1992), 350–373.

239. Micchelli, C. A., Interpolation of scattered data: distance matrices and conditionally positive definite functions, Constr. Approx. 2 (1986), 11–22.

240. Milroy, M. J., G. W. Vickers, and C. Bradley, An adaptive radial basis function approach to modeling scattered data, xxx, to appear.

241. Montès, P., Local kriging interpolation: Application to scattered data on the sphere, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 325–329.

242. Montès, P., Smoothing noisy data by kriging with nugget effects, in Wavelets, Images, and Surface Fitting, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), A. K. Peters, Wellesley MA, 1994, 371–378.

243. Moridis, G. J. and E. J. Kansa, The method of Laplace transform multiquadrics for the solution of the groundwater flow equation, in Advances in Computer Methods for Partial Differential Equations – VII, R. Vichnevetsky, D. Knight, and G. Richter (eds), IMACS, New Brunswick, 1992, xxx–xxx.

244. Moridis, G. J. and E. J. Kansa, The Laplace transform multiquadric method: A highly accurate scheme for the numerical solution of linear partial differential equations, J. Appl. Sc. Comp. 1 (1994), 375–407.

245. Mulgrew, B., Applying radial basis functions, IEEE Signal Proc. Magazine 13 (1996), 50–xx.

246. Muttiah, R., M. Hutchinson, and P. Dyke, Climate surfaces of Texas using thin plate splines, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 427–434.

247. Myers, D. E., Interpolation with positive definite functions, Sciences de la Terre 28 (1988), 252–265.

248. Myers, D. E., Kriging, co-Kriging, radial basis functions and the role of positive definiteness, Comput. Math. Appl. 24 (1992), 139–148.

249. Myers, D. E., Selection of a radial basis function for interpolation, in Advances in Computer Methods for Partial Differential Equations – VII, R. Vichnevetsky, D. Knight, and G. Richter (eds), IMACS, New Brunswick, 1993, xxx–xxx.

250. Narcowich, F. J., Generalized Hermite interpolation and positive definite kernels on a Riemannian manifold, J. Math. Anal. Appl. 190 (1995), 165–193.

251. Narcowich, F. J., N. Sivakumar, and J. D. Ward, On condition numbers associated with radial-function interpolation, J. Math. Anal. Appl. 186 (1994), 457–485.

252. Narcowich, F. J., P. Smith, and J. D. Ward, Density of translates of radial functions on compact sets, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 435–442.

253. Narcowich, F. J. and J. D. Ward, Norms of inverses and condition numbers for matrices associated with scattered data, J. Approx. Theory 64 (1991), 69–94.

254. Narcowich, F. J. and J. D. Ward, Norms of inverses for matrices associated with scattered data, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 341–348.

255. Narcowich, F. J. and J. D. Ward, Norm estimates for the inverses of a general class of scattered-data radial-function interpolation matrices, J. Approx. Theory 69 (1992), 84–109.

256. Narcowich, F. J. and J. D. Ward, Generalized Hermite interpolation via matrixvalued conditionally positive definite functions, Math. Comp. 63 (1994), 661– 687.

257. Narcowich, F. J. and J. D. Ward, Wavelets associated with periodic basis functions, CAT #340, Texas A&M University, 1994.

258. Narcowich, F. J. and J. D. Ward, Nonstationary wavelets on the m-sphere for scattered data, Texas A&M University, 1995.

259. Narcowich, F. J. and J. D. Ward, Nonstationary spherical wavelets for scattered data, in Approximation Theory VIII, Vol. 2: Wavelets and Multilevel Approximation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 301–308.

260. Nielson, G. M., Coordinate free scattered data interpolation, in Topics in Multivariate Approximation, C. K. Chui, L. L. Schumaker, and F. Utreras (eds), Academic Press, New York, 1987, xxx–xxx.

261. Nielson, G. M. and T. A. Foley, A survey of applications of an affine invariant norm, in Mathematical Methods in Computer Aided Geometric Design, T. Lyche and L. Schumaker (eds), Academic Press, New York, 1989, 445–467.

262. Nuss, W. A., Three dimensional meteorological analysis using multiquadric interpolation, Int. J. Sci. Comp. Model., to appear.

263. Nuss, W. A. and D. W. Titley, use of multiquadric interpoaltion for meteorological objective analysis, Monthly Weather Review 12 (1994), 1011–1031.

264. Orr, M. J. L., Regularization in the selection of radial basis function centres, Neural Computation 7 (1995), 606–623.

265. Park, J. and I. W. Sandberg, Universal approximation using radial bais function networks, Neural Computation 3 (1991), 246–257.

266. Park, J. and I. W. Sandberg, Approximation and radial basis function networks, Neural Computation 5 (1993), 305–317.

267. Pegram, C. and G. S. Pegram, Integration of rainfall via multiquadric surfaces over polygons, J. Hydraulic Eng. 114 (1993), 151–163.

268. Poggio, T. and F. Girosi, Networks for approximation and learning, Proc. IEEE 78 (1990), 1481–1497.

269. Pottmann, M. and H. P. Jörgl, Radial basis function networks for internal model control, Appl. Math. Comput. 70 (1995), 283–298.

270. Pottmann, H. and M. Eck, Modified multiquadric methods for scattered data interpolation over a sphere, Comput. Aided Geom. Design 7 (1990), 313–321.

271. Powell, M. J. D., Radial basis functions for multivariable interpolation: a review, in Algorithms for the Approximation of Functions and Data, J. C. Mason and M. G. Cox (eds), Oxford Univ. Press, Oxford, 1987, 143–167.

272. Powell, M. J. D., Radial basis function approximations to polynomials, in Numerical Analysis 1987, D. F. Griffiths and G. A. Watson (eds), Longman Scientific and Technical, Essex, 1987, 223–241.

273. Powell, M. J. D., Univariate multiquadric approximation: reproduction of linear polynomials, in Multivariate Approximation and Interpolation, ISNM 94, W. Haussman and K. Jetter (eds), Birkhäuser, Basel, 1990, 227–240.

274. Powell, M. J. D., Univariate multivariate interpolation: some recent results, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 371–382.

275. Powell, M. J. D., The theory of radial basis functions in 1990, in Advances in Numerical Analysis II: Wavelets, Subdivision, and Radial Basis Functions, W. Light (ed), Oxford University Press, Oxford, 1992, 105–210.

276. Powell, M. J. D., Tabulation of thin-plate splines on a very fine two-dimensional grid, in Numerical Methods in Approximation Theory, ISNM 105, D. Braess, L.L. Schumaker (ed), Birkhäuser, Basel, 1992, 221–244.

277. Powell, M. J. D., Truncated Laurent expansions for the fast evaluation of thin plate splines, in Algorithms for Approximation III, J. C. Mason and M. G. Cox (eds), Chapman and Hall, London, 1992, 99–120.

278. Powell, M. J. D., Some algorithms for thin plate spline interpolation to functions of two variables, Ser. Approx. Decompos. 4 (1994), 303–319.

279. Powell, M. J. D., The uniform convergence of thin plate spline interpolation in two dimensions, Numer. Math. 68 (1994), 107–128.

280. Quak, E., N. Sivakumar, and J. D. Ward, Least squares approximation by radial functions, SIAM J. Numer. Anal. 24 (1993), 1043-1066.

281. Rabut, C., Fast quasi-interpolation of surfaces with generalized B-splines on regular nets, in Mathematics of Surfaces III, D. C. Handscomb (ed), Clarendon Press, Oxford, 1989, 429–449.

282. Rabut, C., B-splines polyharmoniques cardinales: Interpolation, quasi-interpo-lation, filtrage, dissertation, University of Toulouse, 1990.

283. Rabut, C., An introduction to Schoenberg’s approximation, Comput. Math. Appl. 24 (1992), 149–175.

284. Rabut, C., Elementary m-harmonic cardinal B-splines, Numer. Algorithms 2 (1992), 39–67.

285. Rabut, C., High level m-harmonic cardinal B-splines, Numer. Algorithms 2 (1992), 68–84.

286. Rippa, S., Interpolation and smoothing of scattered data by radial basis functions, M.S. Thesis, Tel Aviv University, 1984.

287. Ron, A., The L_{2}-approximation orders of principal shift-invariant spaces generated by a radial basis function, in Numerical Methods in Approximation Theory, ISNM 105, D. Braess, L.L. Schumaker (ed), Birkhäuser, Basel, 1992, 245–268.

288. Ron, A. and X. Sun, Strictly positive definite functions on spheres, Math. Comp., to appear.

289. Saha, A. and C. L. Wu, Approximation, dimension reduction and nonconvex optimization using linear superpositions of Gaussians, IEEE Trans. on Computers 42 (1993), 1222–1233.

290. Šalkauskas, K., Moving least squares interpolation with thin-plate splines and radial basis functions, Comput. Math. Appl. 24 (1992), 177–185.

291. Šalkauskas, K. and L. Bos, Weighted splines as optimal interpolants, Rocky Mountain J. Math. 22 (1992), 205–217.

292. Sanner, R. M. and J. J. E. Slotine, Gaussian networks for direct adaptive control, IEEE Trans. Neural Networks 3 (1992), 837–863.

293. Saundersen, H. C., Multiquadric interpolation of fluid speeds in a natural river channel, Comput. Math. Appl. 24 (1992), 187–193.

294. Schaback, R., Comparison of radial basis function interpolants, in Multivariate Approximation: From CAGD to Wavelets, Kurt Jetter and Florencio Utreras (eds), World Scientific Publishing, Singapore, 1993, 293–305.

295. Schaback, R., Approximation of polynomials by radial basis functions, in Wavelets, Images, and Surface Fitting, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), A. K. Peters, Wellesley MA, 1994, 459–466.

296. Schaback, R., Lower bounds for norms of inverses of interpolation matrices for radial basis functions, J. Approx. Theory 79 (1994), 287–306.

297. Schaback, R., Creating surfaces from scattered data using radial basis functions, in Mathematical Methods for Curves and Surfaces, M. Dæhlen, T. Lyche, and L. Schumaker (eds), Vanderbilt University Press, Nashville, 1995, 477–496.

298. Schaback, R., Error estimates and condition numbers for radial basis function interpolation, Adv. in Comput. Math. 3 (1995), 251–264.

299. Schaback, R., Multivariate interpolation and approximation by translates of a basis function, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 491–514.

300. Schaback, R., Approximation by radial basis functions with finitely many centers, Constr. Approx. 12 (1996), 331–340.

301. Schagen, I. P., Sequential exploration of unknown multidimensional functions as an aid to optimization, IMA J. Numer. Anal. 4 (1984), 337–347.

302. Schimming, R. and M. Belger, Polyharmonic radial functions on a Riemannian manifold, Math. Nachr. 153 (1991), 207–216.

303. Schiro, R. and G. Williams, An adaptive application of multiquadric interpolants for numerically modeling large numbers of irregularly spaced hydrographic data, Survg. Mapp. 44 (1984), 365–381.

304. Schoenberg, I. J., On certain metric spaces arising from Euclidean spaces by a change of metric and their imbedding in Hilbert space, Ann. of Math. 38 (1937), 787–793.

305. Schoenberg, I. J., Metric spaces and completely monotone functions, Ann. of Math. 39 (1938), 811–841.

306. Schoenberg, I. J., Metric spaces and positive definite functions, Trans. Amer. Math. Soc. 44 (1938), 522–536.

307. Schoenberg, I. J., Positive definite functions on spheres, Duke Math. J. 9 (1942), 96–108.

308. Schreiner, M., On a new condition for strictly positive definite functions on spheres, Proc. Amer. Math. Soc., to appear.

309. Sharan, M., E. J. Kansa, and S. Gupta, Application of the multiquadric method for numerical solution of elliptic partial differential equations, Appl. Math. Comp., to appear.

310. Shaw, E. M. and P. P. Lynn, Area rainfall evaluation using two surface fitting techniques, Bull. Int. Ass. Hydrol. Sci. 17 (1972), 419–433.

311. Sibson, R. and G. Stone, Computation of thin-plate splines, SIAM J. Sci. Statist. Comput. 12 (1991), 1304–1313.

312. Sivakumar, N. and J. D. Ward, On the least squares fit by radial functions to multidimensional scattered data, Numer. Math. 65 (1993), 219–243.

313. Sonar, T., Optimal recovery using thin plate splines in finite volume methods for the numerical solution of hyperbolic conservation laws, Universität Göttingen, 1995.

314. Stead, S., Estimation of gradients from scattered data, Rocky Mountain J. Math. 14 (1984), 265–279.

315. Steele, N. C., C. R. Reeves, M. Nicholas, and P. J. King, Radial basis function artificial neural networks for the inference process in fuzzy logic based control, Computing 54 (1995), 99–117.

316. Stewart, J., Positive definite functions and generalizations, an historical survey, Rocky Mountain J. Math. 6 (1976), 409–434.

317. Sun, X., On the solvability of radial function interpolation, in Approximation Theory VI, C. Chui, L. Schumaker, and J. Ward (eds), Academic Press, New York, 1989, 643–646.

318. Sun, X., Multivariate interpolation using ridge or related functions, dissertation, University of Texas, 1990.

319. Sun, X., Norm estimates for inverses of Euclidean distance matrices, J. Approx. Theory 70 (1992), 339–347.

320. Sun, X., Cardinal and scattered-cardinal interpolation by functions having non-compact support, Comput. Math. Appl. 24 (1992), 195–200.

321. Sun, X., Solvability of multivariate interpolation by radial or related functions, J. Approx. Theory 72 (1993), 252–267.

322. Sun, X., Conditionally positive definite functions and their application to multivariate interpolation, J. Approx. Theory 74 (1993), 159–180.

323. Sun, X., Scattered Hermite interpolation using radial basis functions, Linear Algebra Appl. 207 (1994), 135–146.

324. Sun, X., The fundamentality of translates of a continuous function on spheres, Numerical Algorithms 8 (1994), 131–134.

325. Sun, X., Cardinal Hermite interpolation using positive definite functions, Numerical Algorithms 7 (1994), 253–268.

326. Sun, X., Conditional positive definiteness and complete monotonicity, in Approximation Theory VIII, Vol. 1: Approximation and Interpolation, C. Chui, and L. Schumaker (eds), World Scientific Publishing, Singapore, 1995, 537–540.

327. Sun, X. and E. W. Cheney, Fundamental sets of continuous functions on spheres, Southwest Missouri State University, 1994.

328. Tan, S., J. Hao, and J. Vandewalle, Multivariable nonlinear system identification by radial basis function neural networks, Arch. Control Sci. 4 (1995), 55–67.

329. Tarwater, A. E., A parameter study of Hardy’s multiquadric method for scattered data interpolation, Lawrence Livermore National Laboratory, TR UCRL-563670, 1985.

330. Utreras, F. I., Positive thin plate splines, Approx. Theory Appl. 1 (1985), 77–108.

331. Utreras, F. I., Multiresolution and pre-wavelets using radial basis functions, in Multivariate Approximation: From CAGD to Wavelets, Kurt Jetter and Florencio Utreras (eds), World Scientific Publishing, Singapore, 1993, 321– 333.

332. Utreras, F. I. and M. L. Vargas, Monotone interpolations of scattered data in IR^{2}, Constr. Approx. 7 (1991), 49–68.

333. Villalobos, M. A. and G. Wahba, Multivariate thin plate spline estimates for the posterior probabilities in the classification problem, Comm. Statist. A— Theory Methods 12 (1983), 1449–1479.

334. Wahba, G., Convergence rate of ”thin plate” smoothing splines when the data are noisy (preliminary report), Springer Lecture Notes in Math., 757, 1979, 233–245.

335. Wahba, G., Spline Interpolation and smoothing on the sphere, SIAM J. Sci. Statist. Comput. 2 (1981), 5–16.

336. Wahba, G., Erratum: Spline interpolation and smoothing on the sphere, SIAM J. Sci. Statist. Comput. 3 (1982), 385–386.

337. Wahba, G., Multivariate thin plate spline smoothing with positivity and other linear inequality constraints, Statistical image processing and graphics (Luray, Va., 1983), Statist.: Textbooks Monographs, 72, Dekker, New York, 1986, 275–289.

338. Wahba, G., Multivariate model building with additive interaction and tensor product thin plate splines, in Curves and Surfaces, P.-J. Laurent, A. Le Méhauté, and L. L. Schumaker (eds), Academic Press, New York, 1991, 491– 504.

339. Wahba, G., Spline Models for Observational Data, CBMS-NSF Regional Conference Series in Applied Mathematics 59, SIAM, Philadelphia, 1990.

340. Wendland, H., Ein Beitrag zur Interpolation mit radialen Basisfunktionen, Diplomarbeit, Universität Göttingen, 1994.

341. Wendland, H., Piecewise polynomial, positive definite and compactly supported radial functions of minimal degree, Adv. in Comput. Math. 4 (1995), 389–396.

342. Weinrich, M., Charakterisierung von Funktionenräumen bei der Interpolationnn mit radialen Basisfunktionen, dissertation, Universität Göttingen, 1994.

343. Williamson, R. E., Multiply monotone functions and their Laplace transform, Duke Math. J. 23 (1956), 189–207.

344. Wu, Z., Die Kriging-Methode zur Lösung mehrdimensionaler Interpolationsprobleme, dissertation, Universität Göttingen, 1986.

345. Wu, Z., A class of convexity-preserving bases for radial basis interpolation (Chinese), Math. Appl. 3 (1990), 33–37.

346. Wu, Z., Hermite-Birkhoff interpolation of scattered data by radial basis functions, Approx. Theory Appl. 8 (1992), 1–10.

347. Wu, Z., A convergence analysis for a class of radial basis interpolations (Chinese), Gaoxiao Yingyong Shuxue Xuebao 6 (1991), 331–336.

348. Wu, Z., On the convergence of the interpolation with radial basis function, Chinese J. Contemp. Math. 14 (1993), 269–277.

349. Wu, Z., Convergence of interpolation by radial basis functions (Chinese), Chinese Ann. Math. Ser. A 14 (1993), 480–486.

350. Wu, Z., Characterization of positive definite radial functions, in Mathematical Methods for Curves and Surfaces, M. Dæhlen, T. Lyche, and L. Schumaker (eds), Vanderbilt University Press, Nashville, 1995, 573–578.

351. Wu, Z., Compactly supported positive definite radial functions, Adv. in Comput. Math. 4 (1995), 283–292.

352. Wu, Z. and R. Schaback, Local error estimates for radial basis function interpolation of scattered data, IMA J. Numer. Anal. 13 (1993), 13–27.

353. Wu, Z. and R. Schaback, Shape preserving properties and convergence of univariate multiquadric quasi-interpolation, Acta Math. Appl. Sinica (English Ser.) 10 (1994), 441–446.

354. Xu, Y. and E. W. Cheney, Interpolation by periodic radial functions, Comput. Math. Appl. 24 (1992), 201–215.

355. Xu, Y. and E. W. Cheney, Strictly positive definite functions on spheres, Proc. Amer. Math. Soc. 116 (1992), 977–981.

356. Xu, Y., W. A. Light, and E. W. Cheney, Constructive methods of approximation by ridge functions and radial functions, Numer. Algorithms 4 (1993), 205–223.

357. Yamada, T. and L. C. Wrobel, Properties of Gaussian radial basis functions in the dual reciprocity boundary element method, Z. Angew. Math. Phys. 44 (1993), 1054–1067.

358. Zeckzer, D., Dreiecks-basierte lokale Scattered Data Interpolation unter Verwendung radialer Basismethoden, Diplomarbeit, Universität Kaiserslautern,

Related Material

359. Alfeld, P., Scattered data interpolation in three or more variables, in Mathematical Methods in Computer Aided Geometric Design, T. Lyche and L. Schumaker (eds), Academic Press, New York, 1989, 1–33.

360. Aronszajn, N., Theory of reproducing kernels, Trans. Amer. Math. Soc. 68 (1950), 337–404.

361. Askey, R., Orthogonal Polynomials and Special Functions, Reg. Conf. Ser. in Appl. Math., SIAM, 1975.

362. Barnhill, R. E. and H. S. Ou, Surfaces defined on surfaces, Comput. Aided Geom. Design 7 (1990), 323–336.

363. Berg, C., J. P. R. Christensen, and P. Ressel, Harmonic Analysis on Semigroups, Springer, Berlin, 1984.

364. Blumenthal, L. M., Distance Geometries, Univ. of Missouri Studies 13 (1938), 142pp.

365. de Boor, C., Approximation order without quasi-interpolants, in Approximation Theory VII, E. W. Cheney, C. Chui, and L. Schumaker (eds), Academic Press, New York, 1992, 1–18.

366. Chui, C. K., Multivariate Splines, Reg. Conf. Ser. in Appl. Math., SIAM, 1988.

367. Chung, K. C. and T. H. Yao, On lattices admitting unique Lagrange interpolations, SIAM J. Numer. Anal. 14 (1977), 735–743.

368. Cuppens, R., Decomposition of Multivariate Probability, Academic Press, New York, 1975.

369. Donoghue, W. F., Monotone Matrix Functions and Analytic Continuation, Springer, Berlin, 1974.

370. Foley, T. A., Interpolation of scattered data on a spherical domain, in Algorithms for Approximation II, M. G. Cox and J. C. Mason (eds), Chapman & Hall, London, 1990, 303–310.

371. Foley, T. A., D. A. Lane, G. M. Nielson, R. Franke, and H. Hagen, Interpolation of scattered data on closed surfaces, Comput. Aided Geom. Design 7 (1990), 303–312.

372. Foley, T. A., D. A. Lane, G. M. Nielson, and R. Ramaraj, Visualizing functions over a sphere, Comp. Graphics and Applics. 10 (1990), 32–40.

373. Gasper, G., Positivity and special functions, in Theory and Application of Special Functions, R. Askey (ed), Academic Press, New York, 1975, 375–433.

374. Gel’fand, I. M. and G. E. Schilow, Verallgemeinerte Funktionen, Vol. 1, Deutscher Verlag der Wissenschaften, Leipzig, 1960.

375. Gel’fand, I. M. and N. Ya. Vilenkin, Generalized Functions Vol. 4, Academic Press, New York, 1964.

376. Golub, G. H. and C. F. Van Loan, Matrix Computations, Johns Hopkins University Press, Baltimore, 1989.

377. Horn, R. A. and C. R. Johnson, Matrix Analysis, Cambridge University Press, Cambridge, 1985.

378. Horn, R. A. and C. R. Johnson, Topics in Matrix Analysis, Cambridge University Press, Cambridge, 1991.

379. Karlin, S., Total Positivity, Stanford University Press, Stanford, 1968.

380. Lukacs, E., Characteristic Functions, Griffin, London, 1970.

381. Lyche, T., Knot Removal for Spline Curves and Surfaces, in Approximation Theory VII, E. W. Cheney, C. Chui, and L. Schumaker (eds), Academic Press, New York, 1992, 207–226.

382. Marcus, M. and H. Minc, Introduction to Linear Algebra, Dover Publications, New York, 1965 (republished 1988).

383. Müller, C., Spherical Harmonics, Springer Lecture Notes in Mathematics, Vol. 17, 1966.

384. Riesz, F. and B. Sz.-Nagy, Functional Analysis, Dover Publications, New York, 1955 (republished 1990).

385. Stein, E. M. and G. Weiss, Introduction to Fourier Analysis in Euclidean Spaces, Princeton University Press, Princeton, 1971.

386. Szegö, G., Orthogonal Polynomials, Amer. Math. Soc. Coll. Publ. Vol. XXIII, Providence, 1959.

387. Wells, J. H. and R. L. Williams, Embeddings and Extensions in Analysis, Springer, Berlin, 1975.

388. Widder, D. V., The Laplace Transform, Princeton University Press, Princeton,